OpenStudy (anonymous):

how do i solve 7+sqrt5 divied sqrt5-1

OpenStudy (anonymous):

rationalize :)

OpenStudy (anonymous):

\[(7+\sqrt{5}/\sqrt{5}-1) * (\sqrt{5}+1/\sqrt{5}+1)\] we can rationalize using a difference of squares to insure there will not be a radical in the denominator in our answer